OpenStudy (hba):

5. If ten resistances each of 10 ohms are connected in parallel then the total resistance is A) 1 ohm B) 5 ohms C) 2 ohms D) 10 ohms

OpenStudy (hba):

I suppose the answer is 10 ohm

OpenStudy (hba):

1/10+1/10+1/10..1/10 =100/10=10 ohm

OpenStudy (hba):

Then Please tell me how to do it :(

OpenStudy (hba):

@hartnn

hartnn (hartnn):

1/10+1/10+1/10..1/10 = 10/10 = 1

hartnn (hartnn):

methos correct, calculation mistake

hartnn (hartnn):

*method

hartnn (hartnn):

then take reciprocal of 1 which will be 1 only.

OpenStudy (hba):

Oh Wrong calculation

OpenStudy (hba):

Thanks alot hartnn :)

hartnn (hartnn):

welcome ^_^